6. Two dice are tossed. Assume that each possible outcome has a probability. Let A be...

Two dice are tossed. Assume that each possible outcome has a probability. Let A be the event that the sum of the faces showing is 6, and let B be the event that the is twice the face showing on the other. Calculate P(AnB). face showing on one die
Answer #1

Here we have:

Two dice are tossed, assume that each possible outcomes has a 1 / 36 probability.

A = { (1,5), (5,1), (2,4), (4,2), (3,3) }

B = { (1,2), (2,1), (2,4), (4,2), (3,6), (6,3) }

There are 36-6 = 30 events in Bc

However: A and Bc will have only be having 3 pairs together.

(1,5), (5,1), (3,3)

so,  

P(ABc) = 3/36 = 1/12